OpenStudy (anonymous):

Solve: 4 − x = 19 A) 15 B) -15 C) 21 D) -21

OpenStudy (bahrom7893):

add four to both sides

OpenStudy (bahrom7893):

then multiply both sides by -1

OpenStudy (anonymous):

Subtract four from both sides

OpenStudy (anonymous):

yes subtract

OpenStudy (anonymous):

4 - x = 19 - x = 15 Multiply both sides by - 1

OpenStudy (anonymous):

-x=19-4 x= -1*(19-4)

OpenStudy (anonymous):

so the answer is 15?

OpenStudy (anonymous):

No, you're solving from x, right?

OpenStudy (anonymous):

*for

OpenStudy (anonymous):

yes,

OpenStudy (anonymous):

okay so would the answer be C) 21

OpenStudy (anonymous):

No

OpenStudy (anonymous):

you had - x = 15, and you're solving for postive x, so multiply both sides by -1 (whatever you do to one side you must do to the other) and you'll get x = -15

OpenStudy (anonymous):

yes that is true

OpenStudy (anonymous):

thank you all for your help